OpenStudy (anonymous):

2^k -15 is square. Find all k (integers)

OpenStudy (klimenkov):

Does that mean \(2^k-15=a^2, a\in \mathbb N\)?

OpenStudy (anonymous):

Looks like a square in integers to me:-)

OpenStudy (anonymous):

OK, naturals if u insist

OpenStudy (anonymous):

Number theory type question.....

OpenStudy (anonymous):

k=5

OpenStudy (anonymous):

32-15 = 17 is not square

hartnn (hartnn):

k=4,6 as of now

OpenStudy (anonymous):

That's two....

hartnn (hartnn):

next one seems a bit far

OpenStudy (anonymous):

"a bit far" ?

OpenStudy (anonymous):

N is a large place:-)

OpenStudy (anonymous):

If u are saying there are no more solutions, saying it is a bit far is insufficient evidence..:-)

OpenStudy (anonymous):

Hint Consider odd and even k separately

OpenStudy (shubhamsrg):

maybe i got somewhere 2^k = m^2 or 2m^2 for 2^k = m^2 m^2 - 15 = k^2 (m+k)(m-k) = 1*3*5 on comparison, m+k = 5 and m-k =3 =>m=4 -->1st solution again on comparison, m+k =15 m-k =1 m=8 -->2nd solution just have to see through 2m^2 now..

OpenStudy (shubhamsrg):

the k which i used on RHS is different from 2^k

OpenStudy (anonymous):

@shubhamsrg That looks nearly right for even k....

ganeshie8 (ganeshie8):

2^k-15 = p^2 2^k-p^2 = 15 lets say k is even, k = 2t 2^2t - p^2 = 15 (2^t + p)(2^t-p) = 15 solving, we get few solutions lets say k is off, k = 2t+1 2^2t+1 - p^2 = 15 (2^2t+1/2 + p)(2^2t+1/2 - p) = 15 solving we may not get any solutions on cursory check.. only k = 4,6 are the solutions

OpenStudy (shubhamsrg):

similar to what i was going to write..

OpenStudy (anonymous):

For the odd case, use the fact that any square divided by 3 leaves a remainder of 0 or 1

ganeshie8 (ganeshie8):

for the even case, k = {4,6} for the odd case : 2.2^2t+1/2 = 16 2.2^2t+1/2 = 8 we need to solve above both

ganeshie8 (ganeshie8):

again we get the same solutions for odd case k = {4, 6}

OpenStudy (anonymous):

K, for the odd you can also say that 2^k has remainder 2 on division by 3 and then so does 2^k -15 which is not 0 or 1 so there are no solutions when k is odd

OpenStudy (anonymous):